Adult tickets for a play cost $17 and child tickets cost $2. If there were 22 people at a performance and the theater collected $179 from ticket sales, how many adults and how many children attended the play?
Answer by stanbon(48568)

Adult tickets for a play cost $17 and child tickets cost $2. If there were 22 people at a performance and the theater collected $179 from ticket sales, how many adults and how many children attended the play?
--------------------
Let # of adult tickets be "x" ; Value of these is 17x dollars.
# of child tickets is "22-x" ; Value of these is 2(22-x) = 44-2x dollars
-----------------------
EQUATION:
value + value = 179 dollars
17x + 44-2x = 179
15x = 135
x = 9 (# of adult tickets sold)
22-x = 13 (# of child tickets sold)
==============
Cheers,
Stan H.